# UWSAM: Segment Anything Model Guided Underwater Instance Segmentation and A Large-scale Benchmark Dataset

### v1

@@ Abstract
With recent breakthroughs in large-scale modeling, the Segment Anything Model (SAM) has demonstrated significant potential in a variety of visual applications. However, due to the lack of underwater domain expertise, SAM and its variants face performance limitations in end-to-end underwater instance segmentation tasks, while their higher computational requirements further hinder their application in underwater scenarios. To address this challenge, we propose a large-scale underwater instance segmentation dataset, UIIS10K, which includes 10,048 images with pixel-level annotations for 10 categories. Then, we introduce UWSAM, an efficient model designed for automatic and accurate segmentation of underwater instances. UWSAM efficiently distills knowledge from the SAM ViT-Huge image encoder into the smaller ViT-Small image encoder via the Mask GAT-based Underwater Knowledge Distillation (MG-UKD) method for effective visual representation learning. Furthermore, we design an End-to-end Underwater Prompt Generator (EUPG) for UWSAM, which automatically generates underwater prompts instead of explicitly providing foreground points or boxes as prompts, thus enabling the network to locate underwater instances accurately for efficient segmentation. Comprehensive experimental results show that our model is effective, achieving significant performance improvements over state-of-the-art methods on multiple underwater instance datasets. Datasets and codes are available at https://github.com/LiamLian0727/UIIS10K.
Index Terms:
Underwater instance segmentation, Segment anything model, Knowledge distillation.
@@ I Introduction
Underwater visual understanding and dense prediction tasks have recently received increasing attention in the computer vision field in terms of exploring and protecting underwater ecosystems [1, 2, 3, 4, 5]. Instance segmentation, a fundamental and crucial vision task, aims to segment all objects in an image and distinguish them from each other, thereby enabling detailed and accurate scene understanding [6]. By providing pixel-level object localization, instance segmentation provides the foundation for finer instance-level perception and a wide range of applications [7]. Underwater Images Instance Segmentation (UIIS) shows great potential for expanding underwater vision applications, such as discovery of marine relics, exploration of marine resources, underwater human-computer interaction, and comprehensive analysis of underwater images [4].
@@
In recent years, large-scale language models (LLMs) such as the Generative Pretraining Transformer (GPT)-4 [8], Language Learning for Adaptive Multitasking Architecture (LLaMA) [9] and Pathways Language Model (PaLM) [10] have sparked a revolution in the field of natural language processing (NLP). These foundational models exhibit excellent generalization capabilities and perform well in numerous open-world language tasks. Inspired by the success of LLMs, visual base models such as Contrastive Language-Image Pre-Training model (CLIP) [11], Segment Anything Model (SAM) [12]  and Segment Anything Model 2 (SAM2) [13] also emerged. The introduction of these foundation models continues to drive researchers’ exploration in the field of computer vision. Notably, SAM and SAM 2 exhibit robust performance in various segmentation tasks with their powerful encoder-decoder transformer framework and large-scale datasets. With fine-tuning or appropriate modifications, it has significant potential in the field of marine science. Despite these advantages, deploying models like SAM on edge computing devices such as unmanned underwater vehicles still suffers from performance bottlenecks. The primary issue stems from the complexity of the model architectures, especially the image encoders (e.g., ViT-H [14] and Hiera-L [10]), which contain 636M and 224M parameters, respectively. Consequently, using these models for instance segmentation requires considerable computational and memory costs, greatly limiting their applicability in underwater environments.
To address this challenge, several recent studies have proposed strategies to avoid incurring huge costs in prompt-based instance segmentation. For example, FastSAM[15] utilizes a real-time CNN-based architecture to reduce computational overhead, while EfficientSAM [16] leverages the well-known Masked Autoencoders (MAE) method [17] and the SAM model to refine a lightweight and efficient image encoder.
Despite these efforts, SAM and its variants have not been able to achieve satisfactory performance in underwater instance segmentation due to the lack of specialized underwater vision knowledge. To address this limitation, we propose the UIIS10K dataset, a comprehensive large-scale underwater instance segmentation dataset containing 10,048 images with pixel-level annotations in 10 categories such as fish, coral, ruins, and human. Our dataset is intended to provide a solid foundation for underwater instance segmentation tasks, facilitating the development of models that are better adapted to underwater environments.
@@
To evaluate the effectiveness of our proposed methods, we conducted extensive experiments on the UIIS10K datasets, USIS10K datasets [4], and UIIS datasets [2], comparing our approach with state-of-the-art SAM-based methods, as well as other underwater and general instance segmentation methods. The results demonstrate that our approach significantly outperforms these existing methods in terms of segmentation accuracy, robustness, and efficiency. Our methods particularly excel in addressing the challenges unique to underwater environments, such as variable lighting, occlusions, and the complex appearance of underwater objects. The superior performance of our approach highlights its potential as a practical solution for underwater exploration and monitoring, providing accurate and efficient segmentation capabilities even under challenging conditions. The main contributions are concluded as follows:
@@ II-A Underwater Instance Segmentation Dataset
where channel $c\in\left\{R,G,B\right\}$, $I_{c}$ is the image captured by the camera, $J_{c}$ is the scene radiance, $B^{\infty}_{c}$ is the veiling light, $\beta_{c}^{D}$ and $\beta^{B}_{c}$ controls are the light attenuation coefficient and backscattering coefficient of the image in RGB channel [20, 21], and $z$ is the scene depth. In addition to quality degradation, marine organisms also display significant diversity in shape, size, and color [22, 23], all of which result in general models trained on terrestrial instance segmentation datasets not being able to achieve satisfactory results on underwater instance segmentation tasks. Therefore, the community has been focusing on annotating underwater segmentation datasets in recent years. The TrashCan dataset [24] focuses on marine environmental protection and provides annotated underwater trash images for training autonomous underwater vehicles (AUVs) detectors. The DeepFish dataset [25] contains in-air fish images from wholesale fish markets for fish species classification and size estimation. The UIIS dataset [2] is the first generalized underwater instance segmentation dataset containing 4628 images in 7 categories with pixel-level annotations, thus encouraging the development of more accurate and context-sensitive segmentation solutions. The USIS10K dataset [4] focuses on another variant of instance segmentation, underwater salient instance segmentation, and is designed to help segmentation models focus on segmenting valuable underwater salient objects. However, the above datasets are either designed for a particular task, or the number of annotations is insufficient. Since large-scale benchmark datasets play an essential role in developing underwater instance segmentation methods in the large model era, we constructed a new challenging dataset, called USIS10K, which covers 10,048 underwater images with fine-grained annotations for 10 categories of underwater instances. Compared to other existing datasets, our dataset is larger and has more diverse categories of objects.
@@ III UIIS10K DATASET
III-A Dataset Collection & Annotation & Splitting
TABLE I: Comparison with the existing dataset of underwater instances. Where Numbers denotes the number of images in each dataset
@@ III-B Dataset Characteristics and Statistics
Figure 2: Essential characteristics of the UIIS10K dataset. (a) Distribution of the number of instances per image in the UIIS10K dataset. (b) The number of instances per category in the UIIS10K dataset. (c) Distribution of image resolutions in the UIIS10K dataset.
@@
Number and Category of Dataset. A large number of underwater images from different scenes and containing different categories is essential to improve the generalization ability and avoid overfitting of the network in complex marine environments. To address this, we extended the UIIS dataset [2] to provide detailed annotations for 10,048 images from different scenes. As shown in Table I, UIIS10K is the largest existing instance segmentation dataset for underwater scenarios. Furthermore, in order to make the UIIS10K dataset applicable to a broader range of downstream tasks, we labeled the instances in UIIS10K according to the 10 categories shown in Table II. The UIIS10K dataset includes categories such as fish, reptiles, artiodactyla, and ollusks, which are the main research components of marine ecological exploration. It also provides pixel-level annotations for corals, plants, and garbage, which are primary targets in marine ecological conservation. In Fig. 2 (b), we also count the number of masks for each category. the UIIS10K dataset contains a total of 41,862 masks, with the fish and coral categories having the highest number of masks, which are the two most common objects in underwater environments. In addition, the dataset contains annotations for human divers, robots, and ruin, which are valuable for training applications involving human-robot-object intelligent cooperation. Detailed category definitions are given in Table II. In addition, the UIIS10K dataset divides all annotations into three parts: category labels, instance masks, and bounding boxes. This provides the possibility of using the UIIS10K dataset for other downstream tasks in underwater scenarios such as object detection and semantic segmentation.
Number of instances in the image. In the UIIS10K dataset, a single image often contains multiple instances. As shown in Fig. 1, most of these images typically have more than one instance. In Fig. 2 (a), we provide statistics on the number of instances in the dataset. Specifically, for the UIIS10K dataset, 22.83% of the scenes have more than 5 instances, 8.45% of the scenes have more than 10 instances, and the image with the most instances contains 81 instances. The increase in the number of instances in underwater images is often due to factors such as fish groups or coral clusters. For example, in the image in the second row, second column of Fig. 1, instances tend to cluster closely together and occlude each other, which presents a greater challenge for accurately segmenting the instance boundaries.
@@
Size of instances. Segmenting too small or too large instances is a common but challenging problem in the field of instance segmentation. In [45], instances are divided into three levels: small ($area<32^{2}$), medium ($32^{2}\leq area<96^{2}$) and large ($area\geq 96^{2}$), where $area$ denotes the number of pixels in an instance. The number of instances of these three levels in the UIIS10K dataset are 15,800, 13,702, and 12,360, respectively, with a ratio of 1.27:1.05:1. This results in a nearly balanced distribution of instance sizes in UIIS10K, which helps the network learn to segment instances of various sizes effectively.
Various image resolutions and image scenarios. As shown in Fig. 2 (c), UIIS10K dataset includes images of various resolutions, ranging from low-resolution images captured by hand-held cameras to medium-resolution images taken by industrial equipment during underwater missions, with a small number of high-resolution images as well. This diverse resolution range is designed to meet the requirements of different tasks. Additionally, the UIIS10K dataset features images from shallow waters (e.g., sixth column, second row of Fig. 1), images with significantly degraded quality (e.g., fourth column, second row of Fig. 1), images with complex backgrounds (e.g., second column, second row of Fig. 1), and images with high saturation or contrast (e.g., fourth column, first row of Fig. 1). With these scenes, UIIS10K can comprehensively assess the generalizability of the network in different underwater scenarios
@@ IV-C Mask GAT-based Underwater Knowledge Distillation
        Reshape $\mathcal{F}^{1}_{s}$ to Graph $\mathcal{G}^{l}$, and compute the first-order neighborhood $\mathcal{N}^{\,l}_{i}$ of each node using Equation 3;
@@
        Reconstruct the graph $\mathcal{G}^{l}$ as $\mathcal{G}^{l}_{\mathrm{R}}$ using Equation 4-5, and reshape it as ^$\hat{\mathcal{F}}^{l}_{s}$;
@@
Then, we randomly masked most of the node features of the graph G, thus attempting to reconstruct the features ^$\hat{\mathcal{F}}^{l}_{s}$ through GAT and align it with the $l$-th layer features $\mathcal{F}^{l}_{t}$ of the SAM Image Encoder, thus forcing the student network to pay attention to the essential information in the image during the distillation process. When using GAT reconstruction, we first calculate the attentional weights between the masked node $\bar{\mathbf{H}}^{l}_{i}$ and node $\bar{\mathbf{H}}^{l}_{j}$ by using the following equation:
@@
Finally, we reshape the reconstructed graph $\mathcal{G}^{l}_{\mathrm{R}}=\{\mathbf{R}^{l}_{i}|1\leq i\leq hw\}$ as ^$\hat{\mathcal{F}}^{l}_{s}$ and calculate its Mean Square Error (MES) with respect to $\mathcal{F}_{t}$ as distillation loss. To ensure the gradual alignment of intermediate features between the student model and the teacher model, we apply MG-UKD at the 3rd, 6th, 9th, and 12th layers of them. Consequently, the final reconstruction function $\mathcal{L}_{MG-UKD}$ can be expressed as:
@@ IV-D End-to-end Underwater Prompt Generator
$\begin{gathered}\mathcal{F}_{roi}=RoIAlign(\mathcal{F}_{img}+\mathcal{F}_{pe},\mathcal{B})\text{,}\\ \mathcal{F}_{prmt}=MLP(Flatten(Conv_{3\times 3}(\mathcal{F}_{roi})))\text{,}\\ \end{gathered}$ (8)

### v4

# Advancing Marine Research: UWSAM Framework and UIIS10K Dataset for Precise Underwater Instance Segmentation

@@ Abstract
With recent breakthroughs in large-scale models, the Segment Anything Model (SAM) has demonstrated significant potential in a variety of visual applications. However, due to the lack of underwater domain expertise, SAM and its variants face performance limitations in end-to-end underwater instance segmentation tasks, while their higher computational requirements further hinder their application in underwater scenarios. To address this challenge, we propose a large-scale underwater instance segmentation dataset, UIIS10K, which includes 10,048 images with pixel-level annotations for 10 categories. Then, we introduce UWSAM, an approach designed for automatic and accurate segmentation of underwater instances. UWSAM efficiently distills knowledge from the SAM ViT-Huge image encoder into the smaller ViT-Small image encoder via the Mask GAT-based Underwater Knowledge Distillation (MG-UKD) method for effective visual representation learning. Furthermore, we design an End-to-end Underwater Prompt Generator (EUPG) for UWSAM, which automatically generates underwater prompts instead of explicitly providing foreground points or boxes as prompts, thus enabling the network to locate underwater instances accurately for efficient segmentation. Comprehensive experimental results indicate that UWSAM-Student achieves higher segmentation accuracy than WaterMask by 1.3, 0.5, and 1.1 mAPs on UIIS10K, UIIS, and USIS10K respectively, while UWSAM-Teacher surpasses the similarly sized USIS-SAM by 4.8, 1.8, and 2.5 mAPs on the same datasets. Datasets and codes are available at https://github.com/LiamLian0727/UIIS10K.
@@ I Introduction
Figure 1: (a) Comparison of the UIIS and UIIS10K datasets in terms of the number of images and instances. UIIS10K significantly expands the dataset scale. (b) Comparison of models in terms of parameters versus mAPs, UWSAM-Teacher and UWSAM-Student achieving strong performance at different scales. Effi. SAM denotes MaskRCNN[1] + SAM[2]. mAPs is mask AP metrics
@@
Underwater visual understanding and dense prediction tasks have recently received increasing attention in the computer vision field in terms of exploring and protecting underwater ecosystems [3, 4, 5, 6, 7]. Instance segmentation, a fundamental and crucial vision task, aims to segment all objects in an image and distinguish them from each other, thereby enabling detailed and accurate scene understanding [1]. By providing pixel-level object localization, instance segmentation provides the foundation for finer instance-level perception and a wide range of applications [8]. Underwater Images Instance Segmentation (UIIS) shows great potential for expanding underwater vision applications, such as discovery of marine relics, exploration of marine resources, underwater human-computer interaction, and comprehensive analysis of underwater images [4]. The UIIS dataset [4] is the first large-scale dataset for underwater instance segmentation that provides pixel-level annotations in different scenes and object classes, and proposes WaterMask, a detail-enhanced boundary-aware framework for underwater visual features, which provides a solid foundation for subsequent research. However, the UIIS dataset contains only 4628 images, limiting its scale for large-model training. In addition, as a CNN-based approach, WaterMask may achieve lower segmentation accuracy than more recent transformer-based foundation segmentation models such as SAM [9].
@@
In recent years, large-scale language models (LLMs) such as the Generative Pretraining Transformer (GPT)-4 [10], Language Learning for Adaptive Multitasking Architecture (LLaMA) [11] and Pathways Language Model (PaLM) [12] have sparked a revolution in the field of natural language processing (NLP). These foundational models exhibit excellent generalization capabilities and perform well in numerous open-world language tasks. Inspired by the success of LLMs, visual base models such as Contrastive Language-Image Pre-Training model (CLIP) [13], Segment Anything Model (SAM) [9]  and Segment Anything Model 2 (SAM2) [14] also emerged. The introduction of these foundation models continues to drive researchers’ exploration in the field of computer vision. Notably, SAM and SAM 2 exhibit robust performance in various segmentation tasks with their powerful encoder-decoder transformer framework and large-scale datasets. With fine-tuning or appropriate modifications, it has significant potential in the field of marine science. Despite these advantages, deploying models like SAM on edge computing devices such as unmanned underwater vehicles still suffers from performance bottlenecks. The primary issue stems from the complexity of the model architectures, especially the image encoders (e.g., ViT-H [15] and Hiera-L [12]), which contain 636M and 224M parameters, respectively. Consequently, directly utilizing these models for instance segmentation requires considerable computational and memory costs, greatly limiting their applicability in underwater environments.
To address this challenge, several recent studies have proposed strategies to avoid incurring huge costs in prompt-based instance segmentation. For example, FastSAM[16] utilizes a real-time CNN-based architecture to reduce computational overhead, while EfficientSAM [2] leverages the well-known Masked Autoencoders (MAE) method [17] and the SAM model to refine a lightweight and efficient image encoder. Despite these efforts, obtaining strong performance with SAM and its variants in underwater instance segmentation typically requires substantial downstream data for post-training [3, 5]. For example, in underwater salient instance segmentation, the USIS-SAM [6] approach utilizes the USIS10K dataset and an adapter-based fine-tuning strategy to achieve competitive results. However, since USIS10K is specifically designed for salient segmentation, it does not provide annotations for all instances present in the scene. To address this limitation and support large-scale post-training and evaluation of models such as SAM for underwater instance segmentation, we introduce the UIIS10K dataset, containing 10,048 underwater images with pixel-level annotations in 10 categories such as fish, coral, ruins, and human. This dataset is intended to support instance segmentation tasks and facilitate the development of models better adapted to underwater environments. As shown in Figure 1(a), the UIIS10K dataset contains 5,420 additional images and 13,452 more pixel-level instance masks than the UIIS dataset, offering a larger resource for training and evaluation.
@@
[…]
The main contributions are concluded as follows:
@@ II-A Underwater Instance Segmentation Dataset
where channel $c\in\left\{R,G,B\right\}$, $I_{c}$ is the image captured by the camera, $J_{c}$ is the scene radiance, $B^{\infty}_{c}$ is the veiling light, $\beta_{c}^{D}$ and $\beta^{B}_{c}$ controls are the light attenuation coefficient and backscattering coefficient of the image in RGB channel [20, 21], and $z$ is the scene depth. In addition to quality degradation, marine organisms also display significant diversity in shape, size, and color [22, 23], all of which result in general models trained on terrestrial instance segmentation datasets not being able to achieve satisfactory results on underwater instance segmentation tasks. Therefore, the community has been focusing on annotating underwater segmentation datasets in recent years. The TrashCan dataset [24] focuses on marine environmental protection and provides annotated underwater trash images for training autonomous underwater vehicles (AUVs) detectors. The DeepFish dataset [25] contains in-air fish images from wholesale fish markets for fish species classification and size estimation. The UIIS dataset [4] is the first generalized underwater instance segmentation dataset containing 4628 images in 7 categories with pixel-level annotations, thus encouraging the development of more accurate and context-sensitive segmentation solutions. The USIS10K dataset [6] focuses on another variant of instance segmentation, underwater salient instance segmentation, and is designed to help segmentation models focus on segmenting valuable underwater salient objects. However, the above datasets are either designed for a particular task, or the number of annotations is insufficient. Since large-scale benchmark datasets play an essential role in developing underwater instance segmentation methods in the large model era, we constructed a new challenging dataset, called USIS10K, which covers 10,048 underwater images with fine-grained annotations for 10 categories of underwater instances. Compared to other existing datasets, our dataset is larger and has more diverse categories of objects. The refUDS dataset [26] extends UIIS and USIS10K by pairing detailed textual descriptions of underwater images with their corresponding segmentation masks. This resource enables the development and training of models capable of jointly processing visual and linguistic inputs.
@@ III UIIS10K DATASET
TABLE I: Comparison with the existing underwater instances dataset. Numbers denotes the number of images in each dataset
@@
III-A Dataset Collection & Annotation & Splitting
Figure 3: The number of instances per category in the UIIS10K dataset. There are 41,862 carefully labeled instance masks in the UIIS10K dataset.
@@ III-A Dataset Collection & Annotation & Splitting
Figure 4: Distribution of the number of instances per image in the UIIS10K dataset.
@@ III-B Dataset Characteristics and Statistics
Number and Category of Dataset. A large number of underwater images from different scenes and containing different categories is essential to improve the generalization ability and avoid overfitting of the network in complex marine environments. To address this, we extended the UIIS dataset [4] to provide detailed annotations for 10,048 images from different scenes. As shown in Table I, UIIS10K is the largest existing instance segmentation dataset for underwater scenarios. Furthermore, in order to make the UIIS10K dataset applicable to a broader range of downstream tasks, we labeled the instances in UIIS10K according to the 10 categories shown in Table II. The UIIS10K dataset includes categories such as fish, reptiles, artiodactyla, and ollusks, which are the main research components of marine ecological exploration. It also provides pixel-level annotations for corals, plants, and garbage, which are primary targets in marine ecological conservation. In Fig. 3, we also count the number of masks for each category. the UIIS10K dataset contains a total of 41,862 masks, with the fish and coral categories having the highest number of masks, which are the two most common objects in underwater environments. In addition, the dataset contains annotations for human divers, robots, and ruin, which are valuable for training applications involving human-robot-object intelligent cooperation. Detailed category definitions are given in Table II. In addition, the UIIS10K dataset divides all annotations into three parts: category labels, instance masks, and bounding boxes. This provides the possibility of using the UIIS10K dataset for other downstream tasks in underwater scenarios such as object detection and semantic segmentation.
Number of instances in the image. In the UIIS10K dataset, a single image often contains multiple instances. As shown in Fig. 2, most of these images typically have more than one instance. In Fig. 4, we provide statistics on the number of instances in the dataset. Specifically, for the UIIS10K dataset, 22.83% of the scenes have more than 5 instances, 8.45% of the scenes have more than 10 instances, and the image with the most instances contains 81 instances. The increase in the number of instances in underwater images is often due to factors such as fish groups or coral clusters. For example, in the image in the second row, second column of Fig. 2, instances tend to cluster closely together and occlude each other, which presents a greater challenge for accurately segmenting the instance boundaries.
@@
Figure 6: Distribution of image resolutions in the UIIS10K dataset.
@@
Size of instances. Segmenting too small or too large instances is a common but challenging problem in the field of instance segmentation. In [46], instances are divided into three levels: small ($area<32^{2}$), medium ($32^{2}\leq area<96^{2}$) and large ($area\geq 96^{2}$), where areaarea denotes the number of pixels in an instance. The number of instances of these three levels in the UIIS10K dataset are 15,800, 13,702, and 12,360, respectively, with a ratio of 1.27:1.05:1. This results in a nearly balanced distribution of instance sizes in UIIS10K, which helps the network learn to segment instances of various sizes effectively.
Various image resolutions and image scenarios. As shown in Fig. 6, UIIS10K dataset includes images of various resolutions, ranging from low-resolution images captured by hand-held cameras to medium-resolution images taken by industrial equipment during underwater missions, with a small number of high-resolution images as well. This diverse resolution range is designed to meet the requirements of different tasks. Additionally, the UIIS10K dataset features images from shallow waters (e.g., sixth column, second row of Fig. 2), images with significantly degraded quality (e.g., fourth column, second row of Fig. 2), images with complex backgrounds (e.g., second column, second row of Fig. 2), and images with high saturation or contrast (e.g., fourth column, first row of Fig. 2). With these scenes, UIIS10K can comprehensively assess the generalizability of the model in different underwater scenarios.
@@ IV-C Mask GAT-based Underwater Knowledge Distillation
   Reshape $\mathcal{F}^{l}_{s}$ to Graph $\mathcal{G}^{l}$, and compute the first-order neighborhood $\mathcal{N}^{\,l}_{i}$ of each node using Equation 3;
@@
   Reconstruct the graph $\mathcal{G}^{l}$ as $\mathcal{G}^{l}_{\mathrm{R}}$ using Equation 4-5, and reshape it as $\hat{\mathcal{F}}^{l}_{s}$;
@@
Then, we randomly masked most of the node features of the graph G, thus attempting to reconstruct the features $\hat{\mathcal{F}}^{l}_{s}$ through GAT and align it with the $l$-th layer features $\mathcal{F}^{l}_{t}$ of the SAM Image Encoder, thus forcing the student network to pay attention to the essential information in the image during the distillation process. When using GAT reconstruction, we first calculate the attentional weights between the masked node $\bar{\mathbf{H}}^{l}_{i}$ and node $\bar{\mathbf{H}}^{l}_{j}$ by using the following equation:
@@
Finally, we reshape the reconstructed graph $\mathcal{G}^{l}_{\mathrm{R}}=\{\mathbf{R}^{l}_{i}|1\leq i\leq hw\}$ as $\hat{\mathcal{F}}^{l}_{s}$ and calculate its Mean Square Error (MES) with respect to $\mathcal{F}_{t}$ as distillation loss. To ensure the gradual alignment of intermediate features between the student model and the teacher model, we apply MG-UKD at the 3rd, 6th, 9th, and 12th layers of them. Consequently, the final reconstruction function $\mathcal{L}_{MG-UKD}$ can be expressed as:
@@ IV-D End-to-end Underwater Prompt Generator
[…]